The length of a rectangle is twice its width. If the area of the rectangle is 200yd^2 , find its perimeter.

Answer by checkley77(12844)

L=2W
LW=200
2W*W=200
2W^2=200
W^2=200/2
W^2=100
W=SQRT100
W=10 ANS. FOR THE WIDTH.
L=2*10=20 ANS. FOR THE LENGTH.
PERIMETER=2L+2W
P=2*20+2*10=40+20=60 YDS. IS THE PERIMETER.